A cake calls for 2 1/2 cups of sugar. The icing calls for 3 3/4 cups of sugar. How much total sugar will you need to make this cake and icing?
Answer by stanbon(75887)

A cake calls for 2 1/2 cups of sugar. The icing calls for 3 3/4 cups of sugar. How much total sugar will you need to make this cake and icing?
-------------
Ans: 2 1/2 + 3 3/4
---
= 5/2 + 15/4
---
= 10/4 + 15/4
----
= 25/4
=============
Cheers,
Stan H.
==========